OpenStudy (anonymous):

American airlines charges $50 for a checked bag weighing over 23 kg but not more than 32 kg. The charge for a checked bag weighing over 32 kg but not more than 45 kg is $100. How much must be paid for three pieces of luggage weighing 27.9 kg, 31.5 kg, and 40.2 kg?

OpenStudy (ganpat):

27.9 kg = 50 (27.93 - 23) = 31.5 kg = (31.5 - 23) = 40.2 kg = ( 40.2 - 32) = add all the three answers, that would be the final answer...

OpenStudy (ganpat):

@cathy.poff : ???

OpenStudy (anonymous):

I'm here, just away for a sec.

OpenStudy (anonymous):

Thanks:-)

OpenStudy (ganpat):

so got the answer ?